OLUTE. U Opis 5 of 13 (10 complete HW Score: 68 13% 27 25 of 40 pts P 13-6 (similar to Question Avicorp has $115 million debt need a and b please
Answer #1

The yield to maturity(YTM) is the annual return from an investment purchased today and held till maturity, i.e., it is the rate at which the current market price of the bond is equal to the present value of all the cash flows from the bond. It is equal to cost of debt.

Let par value be $100

Coupon rate = 6.1%

Time to maturity(t) = 5 years

No of periods = 5 * 2 = 10

Coupon payment = 6.1% of 100 / 2 = $3.05

Present value(PV) = $96 (96% of par)

Ytm/cost of debt =?

Using a financial calculator we get ytm as = 3.53%

Using excel,

Present Value 96
pmt 3.05
no periods 10
Future Value 100
YTM 3.53% (=Rate(10,3.05,-96,100,0)

Ytm(semi annual) = 3.53

Ytm annual = 3.53 * 2 = 7.06%

a) Hence cost of debt = 7.0635%

b) After-tax cost of debt = Cost of debt * (1 - Tax rate)

Tax rate = 40%

After tax cost of debt = 7.06 * (1 - 0.40) = 4.236%

Hence after tax cost of debt is 4.236%.

If you have any doubts please let me know in the comments. Please give a positive rating if the answer is helpful to you. Thanks.